If the gravitationl force between two objects was proportional to 1/R,where R is the separation between them,then a particle in circular orbit under such a force would have its orbital speed V proportional to

Grade:10

1 Answers

Sanju
6 years ago
Given that Grav Force is proportional to 1/R..
So,  Grav Force = GMm/R
We have, centripetal force = mv^2 / R
 
For particle revolving in a circular orbit,  Grav Force = centripetal force
 
Therefore, GMm/R = mv^2 / R
 
or v^2 = GM.
 
So the orbital speed v is proportional to R^0